### v5

# BaCo2(AsO4)2: Strong Kitaev, After All

Abstract
The inelastic neutron scattering results and their analysis unequivocally point to a dominant Kitaev interaction in the honeycomb-lattice cobaltate BaCo2(AsO4)2. Our anisotropic-exchange model closely describes all available neutron scattering data in the material’s field-polarized phase. The density-matrix renormalization group results for our model are in close accord with the unusual double-zigzag magnetic order and the low in-plane saturation field of BaCo2(AsO4)2.
BaCo2(AsO4)2, hereafter BaCAO—a member of the family of layered honeycomb-lattice transition-metal compounds [1] originally studied as possible realizations of the two-dimensional (2D) $XY$ model [2, 3, 4, 5, 6, 7, 8, 9]—has been a mystery: see Ref. [10] for an earlier review. Its magnetic order remains discrepant, with the classical modeling yielding a spiral state but failing to describe its excitation spectrum [8, 9, 10, 11], and the more recent neutron-polarization analysis proposing an unusual double-zigzag state instead [12, 13]. Equally perplexing has been the exceptionally low in-plane critical field to a magnetically-polarized state [4, 10], suggesting that BaCAO resides on the brink of a quantum phase transition.
This enigmatic track record has recently propelled BaCAO into being considered a candidate for proximity to an exotic spin-liquid state [14], as the search for a realization of the Kitaev model [15] has shifted from compounds of $5d$ Ir and $4d$ Ru elements [16, 17, 18, 19] to $3d$ cobaltates [20, 21, 22, 23, 24]. For $d^{7}$ Co2+ ions in an edge-sharing octahedral environment, spin-orbit coupling yields $j_{\text{eff}}\!=\!1/2$ moments and can lead to strong Kitaev exchange via additional active orbitals and hopping paths [22, 23, 24]. However, most Co2+ compounds exhibit ferromagnetic (FM) or zigzag (ZZ) order , with BaCAO being a stark exception [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
Recent studies on many aspects of BaCAO, such as spin excitations, phase diagram, thermal transport, and magnetoelastic properties [37, 38, 39, 40, 14, 41, 42, 43, 44, 45], have all confirmed its status as an unusual material, but have not provided clear resolution on its model [24]. This is despite the smallness of the in-plane critical fields that make the fully polarized FM state easily accessible [10, 14], allowing one to extract BaCAO’s model parameters from the inelastic neutron scattering (INS) spin-flip spectra [46, 47, 48, 49, 50, 51, 52, 53].
Earlier such attempts that relied on the easy-plane ${\sf J}_{1}$${\sf J}_{2}$${\sf J}_{3}$ XXZXXZ model led to a conundrum: the extracted parameters that fit the high-field INS data did not correspond to the expected spiral ground state, nor was the model able to explain the observed fully gapped zero-field spectrum even when assuming such a state [9, 10].
A recent study provided more INS data [40], but rejected substantial Kitaev terms in BaCAO in favor of largely the same easy-plane ${\sf J}_{1}$${\sf J}_{3}$ XXZXXZ model, imposing the zero-field order as a classical spiral state [11]. This model yielded only a rough fit to a subset of the high-field INS data and could not reconcile puzzling low critical fields. The growing enigma became even more bewildering as recent theoretical works suggested a complete absence of the spiral phase in the phase diagram of the quantum S=1/2S\!=\!1/2 ${\sf J}_{1}$${\sf J}_{3}$ XXZXXZ model [54, 55].
However, the easy-plane ${\sf J}_{1}$${\sf J}_{3}$ description of BaCAO has recently received nearly unanimous  [^1] One exception is Ref. [66] by one of the present authors, which uses the same easy-plane $XXZ$ description of the nearest-neighbor interactions, but introduces strong bond-dependent terms in the third-neighbor exchange. support from several first-principles approaches [57, 58, 59, 42]. Thus, despite its inadequate account of BaCAO’s phenomenology, the view of BaCAO as a more ordinary easy-plane ferro-antiferromagnet seemed to be prevailing. Until now.
In this Letter, we present INS data together with analytical and numerical results, which unequivocally point to a dominant Kitaev term in BaCAO’s model. This model closely describes the entire corpus of available INS data for BaCAO’s spectrum above the critical field, demonstrates its proximity to a phase boundary, identifies BaCAO’s ground state as the non-classical double-zigzag (dZZ) state with an out-of-plane tilt of spins in accord with neutron polarimetry [12, 13], and confirms BaCAO’s low in-plane critical fields.
Figure 1: Magnon energy vs $\bf q$} for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$; B=3B\!=\!3 T. Diamonds [9], squares [40], and circles (this work) are the INS data, and solid lines are theory fits. $\Delta E_{\rm M}$ is the energy splitting at M1. Insets: Crystallographic axes and principal directions, 1st-, 2nd-, and 3rd-neighbor bonds on the honeycomb lattice with bond-dependent exchanges represented by colors, and the Brillouin zone with $\Gamma$ and M1 points indicated.
Key argument.—While truly indispensable for inferring magnetic couplings, the spectra in the polarized phase of Kitaev magnets can still be challenging to analyze. Specifically, the spin-flip energies along the high-symmetry momentum directions are often independent of the very bond-dependent terms that these spectra are expected to identify [53, 60, 61]. For data sets dominated by such directions, it can be tempting to conclude that such terms are simply absent from the model.
[…]
$\mathcal{H}\!=\!\sum_{n,\langle ij\rangle_{n}}\!{\sf J}_{n}\Big{(}S^{x}_{i}S^{x}_{j}\!+\!S^{y}_{i}S^{y}_{j}\!+\!\Delta_{n}S^{z}_{i}S^{z}_{j}\Big{)}\!-\!\sum_{i}\mathbf{H}\!\cdot\!\mathbf{\hat{g}}\!\cdot\!\mathbf{S}_{i},$ (1)
where $\langle ij\rangle_{n}$ for n=1,2,3n\!=\!1,2,3 are the $n$th-neighbor bonds, ${\sf J}_{n}$ are the exchanges (${\sf J}_{1,2}\!<\!0$, ${\sf J}_{3}\!>\!0$), and $\Delta_{n}$ are the $XXZ$ anisotropy parameters for such bonds. In BaCAO, ${\sf J}_{2}$ is weaker than ${\sf J}_{3}$, as is typical in the honeycomb-lattice magnets [10, 17], and was neglected in Ref. [40]. All $\Delta_{n}$ are easy-plane-like, ${\bf H}\!=\!\mu_{B}{\bf B}$, and the g-tensor is diagonal, $\mathbf{\hat{g}}\!=\!\{g_{x},g_{y},g_{z}\}$, with $g_{x}\!\approx\!g_{y}\!\approx\!5.0$ and $g_{z}\!\approx\!2.5$ [10], where we use crystallographic frame in which $x\!=\!a^{*}$ and y=by\!=\!b; see inset of Fig. 1.
Because of the $U(1)$ symmetry of the $XXZ$ model (1), one does not expect any dependence of the spectra in the polarized phase on the in-plane field orientation.
[…]
The demonstrated energy splitting $\Delta E_{\rm M}$ at the M1 point in Fig. 1 is about half of the magnon’s bandwidth, suggesting that the Kitaev-like exchanges should be on par with the $XXZ$ terms of model (1). The solid lines use the model discussed below, which will make the magnitude of these terms precise and explicit.
Figure 2: (a) The field dependence of $E_{\Gamma}$ (circles) and $E_{\rm M_{1}}$ (diamonds) for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$, shown as black and blue symbols (data) and lines (theory), respectively. $\Delta E_{\rm M}$ is indicated and shown in the inset. (b) and (c), Same as in Fig. 1: data and their fits for magnon energies vs $\bf q$} with ${\bf B}\!\parallel\!b$, B=0.7B\!=\!0.7 T, 3.0 T, and 6.0 T from Refs. [9, 10] in (b), and with ${\bf B}\!\parallel\!a^{*}$, B=0.6B\!=\!0.6 T, 1.0 T, 2.0 T, and 3.0 T (this work) in (c). Dashed lines represent the lowest edges of the two-magnon continua for 0.7 T in (b), 0.6 T in (c), and for the $\rm M_{1}$} point in (a).
The large band splitting for the two field orientations is further exposed in Figure 2(a), which shows the field dependence of the energies at the $\Gamma$ and M1 points. While $E_{\Gamma}$ data (circles) are nearly degenerate for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$ (black and blue color, respectively), indicating $g_{b}\!\approx\!g_{a^{*}}$, $E_{\rm M_{1}}$ results (diamonds) exhibit a large gap $\Delta E_{\rm M}$ between the two orientations, which is maximal at lower fields. Solid lines show theoretical fits of the $E_{\Gamma(\rm M_{1})}$ data, and the inset shows the fit of the gap $\Delta E_{\rm M}$ vs field.
Figure 3: (a) Data from Ref. [40] for ${\bf B}\!\parallel\!b$, B=3B\!=\!3 T, lines are as described in the legend. (b) Intensity maps of ${\cal S}({\bf q},\omega)$ from Ref. [40] and our results. Dashed lines mark the bottom of the two-magnon continuum. (c) The ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)++(2) with the LT and DMRG boundaries and phases; $\Delta_{n}$, ${\sf J}_{2}/{\sf J}_{1}$, and ${\sf J}_{\pm\pm}/{\sf J}_{1}$ are from (3), symbols show various parameter sets (see the text). (d) Results from DMRG $12\!\times\!12$ cluster calculations for the set (3) with spins in the $yz$ plane.
[…]
We also highlight the two-magnon continua, with their lowest edges shown by dashed lines in Figs. 2(b) and 2(c) for the fields of 0.7 T and 0.6 T, respectively. These lines demonstrate the overlap of the continua with the single-magnon branches that may be responsible for the deviations of the observed energies from the theoretical fits. Dashed lines in Fig. 2(a) serve the same purpose.
Model parameters.—The symmetry-allowed bond-dependent terms in the 1st-neighbor exchange matrix of BaCAO [24], missing from the $XXZ$ model (1), are
${\cal H}_{bd}\!=\!\sum_{\langle ij\rangle_{1}}\!\Big{\{}{\sf J}_{z\pm}\Big{(}\big{(}S^{x}_{i}S^{z}_{j}\!+\!S^{z}_{i}S^{x}_{j}\big{)}\tilde{s}_{\alpha}\!-\!\big{(}S^{y}_{i}S^{z}_{j}\!+\!S^{z}_{i}S^{y}_{j}\big{)}\tilde{c}_{\alpha}\Big{)}$
$\ \ \ +\!2{\sf J}_{\pm\pm}\Big{(}\big{(}S^{x}_{i}S^{x}_{j}\!-\!S^{y}_{i}S^{y}_{j}\big{)}\tilde{c}_{\alpha}\!-\!\big{(}S^{x}_{i}S^{y}_{j}\!+\!S^{y}_{i}S^{x}_{j}\big{)}\tilde{s}_{\alpha}\Big{)}\Big{\}},$ (2)
where $\tilde{c}(\tilde{s})_{\alpha}\!=\!\cos(\sin)\tilde{\varphi}_{\alpha}$ and $\tilde{\varphi}_{\alpha}\!=\!\{0,2\pi/3,-2\pi/3\}$ are the bond angles with the $x$ axis in Fig. 1. The Kitaev exchange is given by $K\!=\!\sqrt{2}{\sf J}_{z\pm}\!-2{\sf J}_{\pm\pm}$. For the full translation of the $XXZ$ and (2) model to the Kitaev notations, $\{{\sf J}_{1},\Delta_{1}{\sf J}_{1},{\sf J}_{\pm\pm},{\sf J}_{z\pm}\}\!\Leftrightarrow\!\{J,K,\Gamma,\Gamma^{\prime}\}$; see EM [19, 17].
[…]
BaCAO’s magnetization data suggest that its field-induced state, from just above the in-plane critical field $H_{c}\!\approx\!0.55$ T, is nearly fully polarized and free from quantum fluctuations  [40, 38, 44, 14]. Therefore, one can extract exchange parameters from the INS spin-flip spectra using spin-wave theory (SWT) for such a state.
Our model contains seven terms: symmetry-allowed nearest-neighbor exchanges, $\{{\sf J}_{1},\Delta_{1},{\sf J}_{\pm\pm},{\sf J}_{z\pm}\}$, third-neighbor $XXZ$ couplings, $\{{\sf J}_{3},\Delta_{3}\}$, and a small ${\sf J}_{2}$ [10], which is kept $XY$-like, $\Delta_{2}\!=\!0$, to avoid overfitting. We also used slightly different in-plane g-factors, $g_{a*}\!=\!4.8$ and $g_{b}\!=\!4.85$, to fit the field-dependencies of the energies at the $\Gamma$ and M points in Fig. 2.
We use the INS data for the reference fields of 3 T and 6 T to fix different combinations of the exchanges by the energies at high-symmetry $\bf q$}-points, such as $\Gamma$ and M-points in Figs. 1 and 2. Even at 3 T, the two-magnon continuum can cross the single-magnon branch and affect the data, as can be seen in Fig. 3(a) for the proximity of the K-point. This makes the use of straightforward quality fit criteria more problematic, because the SWT K-point energy needs to be higher than the data.
Delegating further technical details of the parameters’ extraction to Supplemental Material (SM) [62], the proposed BaCAO parameters (in meV, except for $\Delta_{n}$) are
$\{{\sf J}_{1},\Delta_{1},{\sf J}_{\pm\pm},{\sf J}_{z\pm}\}=\{-6.54,0.36,0.15,-3.76\}$ (3)
and $\{{\sf J}_{2},{\sf J}_{3},\Delta_{3}\}\!=\!\{-0.21,1.70,0.03\}$.
This is the set of parameters that provides a close fit to all available INS data for BaCAO in the polarized phase, as shown in Figs. 1 and 2, and Figs. 3(a) and 3(b) for ${\bf B}\!\parallel\!b$ at B=3B\!=\!3 T with the data from Ref. [40]. In Fig. 3(a), we show the data in multiple $\bf q$}-panels together with the $XXZ$ fit from Ref. [40] (black dashed lines) and our results for the single-magnon branches and the bottom of the two-magnon continuum (solid and light-blue dashed lines, respectively). Fig. 3(b) shows a side-by-side comparison of the intensity maps of the dynamical structure factor ${\cal S}({\bf q},\omega)$ from Ref. [40] and our theoretical results. The comparison of our results with the data for B=0.75B\!=\!0.75 T and ${\bf B}\!\parallel\!b$ from Ref. [40] is shown in EM.
We note that the $XXZ$ sector of our BaCAO parameters in (3) is close to those in earlier studies, Refs. [9, 10], with $\Delta_{1}\!\approx\!0.4$, ${\sf J}_{3}/|{\sf J}_{1}|\!\approx\!0.26$, and ${\sf J}_{2}/{\sf J}_{1}\!\approx\!0.034$ compared to our $\Delta_{1}\!=\!0.36$, ${\sf J}_{3}/|{\sf J}_{1}|\!=\!0.26$, and ${\sf J}_{2}/{\sf J}_{1}\!=\!0.032$  [^2] The absolute values of exchanges in Refs. [9, 10] used an additional factor of 2 in their conventions. Since they describe the earlier INS data well [62], it suggests that $\sf J$$_{\pm\pm}$ in (2) should be small, as is the case in our proposed set (3). A large $\sf J$$_{\pm\pm}$ would also yield a field-dependence of energies for ${\bf B}\!\parallel\!a^{*}$ that is incompatible with the new INS data [62]. This leaves a necessarily large ${\sf J}_{z\pm}$ as the only term capable of explaining the $\Delta E_{M}$ gap in Figs. 1 and 2(a), which is our key result.
Last, but not least, translating the nearest-neighbor exchanges in (3) to the Kitaev notations yields
$\{J,K,\Gamma,\Gamma^{\prime}\}\approx\{-3.3,-5.6,3.0,0.6\}\text{ meV},$ (4)
with a dominant Kitaev term, approximately twice the magnitude of either $J$ or $\Gamma$.
Phase diagram.—Our parameters for BaCAO (3) also resolve the enigma of its ground state and identify its place in a broader phase diagram of Kitaev magnets.
In Fig. 3(c), we present the classical Luttinger-Tisza (LT) ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)++(2) for $\Delta_{1}\!=\!0.36$ and ${\sf J}_{2}/{\sf J}_{1}\!=\!0.03$ as in (3), and we set $\Delta_{3}\!=\!{\sf J}_{\pm\pm}\!=\!0$ for simplicity [19, 64]. The sequence of the FM, ZZ, and helical phases shown in Fig. 3(c) is common in many classical honeycomb-lattice models [11, 54, 19].
The red square and white circle represent projections of the $XXZ$ models of BaCAO from Refs. [9, 10] and Ref. [40], respectively, with the first in the FM phase and the second matching the observed ordering vector ${\bf k}\!\approx\!(0.27,0)$ r.l.u. as a classical spiral [40]. Our proposed set (3) is represented by the red diamond. Classically, it also corresponds to a spiral with ${\bf k}\!\approx\!(0.16,0)$ r.l.u. and small in-plane critical fields $\lesssim\!0.15$ T. The green diamond represents a set with similar classical characteristics and provides slightly better fits to the ${\bf B}\!\parallel\!a^{*}$ INS spectra in Fig. 2 [62]. The empty diamond represents a set that is also similar classically, showing comparable, but slightly worse overall quality of fit to the INS data; see EM.
Recent studies of the quantum S=1/2S\!=\!1/2 edition of the XXZXXZ ${\sf J}_{1}$${\sf J}_{3}$ model using the density-matrix renormalization group (DMRG) method [54, 55] found no trace of the spiral phase in its phase diagram, uncovering instead only narrow ranges of non-classical states, such as dZZ. Here, we extend this study to the ${\sf J}_{1}$${\sf J}_{z\pm}$${\sf J}_{3}$ model, with the remaining parameters ($\Delta_{n}$, $\sf J$$_{\pm\pm}$, and ${\sf J}_{2}$) fixed according to the proposed BaCAO set in (3).
In Fig. 3(c), we show the boundaries of the intermediate phases, which are inferred by extrapolation from several DMRG “scans,” shown by horizontal and vertical arrows. Many individual points were also checked by DMRG (“non-scans”) to verify the phases. DMRG calculations [65], which used the same approach for a closely related model, have been performed recently [19] with a similar goal of exploring intermediate phases between FM and ZZ; see EM and Ref. [19] for technical details.
Similar to prior works [54, 55], the spiral phase is eliminated along the $XXZ$ line. However, at finite ${\sf J}_{z\pm}$, a sequence of non-classical quasi-collinear phases emerges. At lower ${\sf J}_{z\pm}$, the triple-zigzag state, not unlike the one briefly discussed in Ref. [55], occurs, while for the region relevant to BaCAO, a dZZ state firmly establishes itself. For larger ${\sf J}_{z\pm}$, we observed a state that resembles a spin liquid (SL), with only short-ranged correlations. However, further analysis of this state is needed. The set highlighted by a green diamond is in the ZZ phase.
DMRG simulation of the proposed parameter set (3) on a $12\times 12$ cylinder are shown in Fig. 3(d), demonstrating the dZZ state with spins shown in the $yz$ plane to emphasize their out-of-plane tilt. While the non-classical dZZ state for the XXZXXZ ${\sf J}_{1}$${\sf J}_{3}$ models has recently received theoretical endorsements [42, 66, 58, 55], the tilt is unusual and is clearly due to the ${\sf J}_{z\pm}$ coupling. According to our DMRG results, every second spin in the dZZ structure is tilted out of plane by about $35\degree$. Using the g-factors $g_{z}/g_{ab}\!\approx\!0.5$, this yields $7.8\degree$ average tilt of the magnetic moment, compared with $\sim\!6\degree$ suggested by neutron polarimetry [12, 13]. These and other quantitative discrepancies can be attributed either to finite-size effects in DMRG or to small terms in the description of BaCAO that are not accounted for by our model [62].
As one can see in Fig. 3(c), the proposed parameter set (3) appears close to the dZZ-ZZ phase boundary within the dZZ phase, likely related to the anomalously low in-plane critical fields in BaCAO, $g_{ab}\mu_{B}B_{c}/|{\sf J}_{1}|\!\simeq\!0.02$ [2]. Using the DMRG scans vs field [62], we have found critical fields for our model of BaCAO slightly higher than the observed 0.55 T values: $B_{c}^{(b)}\!\approx\!B_{c}^{(a^{*})}\!\approx\!0.7$ T, but nearly equal, in agreement with Refs. [42, 43]. The out-of-plane critical field obtained by the DMRG scan is about 47 T [62], consistent with experiments [42]. Lastly, the field-induced “up-up-down” columnar state—present in the BaCAO’s phase diagram below the polarized state—has also been captured by our model (see SM [62]).
Summary.—Our study brings the multifaceted conundrum of BaCAO close to a complete resolution. The INS data presented in this work provide the most direct evidence of a dominant Kitaev term in its model, establishing BaCAO as a Kitaev-champion among cobaltates and comparable to, if not superior to, other Kitaev magnets, all while reconciling its puzzling ground state and low critical fields with its model’s phenomenology.
Our nearly ideal fit of all available INS spectra in the polarized state also includes nuances such as analyses of the crossings with the two-magnon continuum. The ground state of our model for BaCAO corresponds to the non-classical double-zigzag state with a finite tilt of spins out of the crystallographic plane, all in accord with the neutron polarimetry data. The proposed parameter set for BaCAO is shown to correspond to near proximity to a phase boundary, with small and nearly isotropic in-plane and large out-of-plane critical fields, all consistent with the material’s phenomenology.
For the broader field of Kitaev magnets, the striking variety of phases present between the ferromagnetic and zigzag states in their models’ phase diagrams calls for further in-depth studies. Inspired by the BaCAO example, explorations of materials that potentially host bond-dependent interactions should be pursued to uncover novel phases and phenomena.
Acknowledgements.
Acknowledgments.—We are greatly indebted to T. Halloran and C. Broholm for sharing their published inelastic neutron scattering data. We also thank S. Streltsov and S. Winter for their useful discussions, and Chaebin Kim for raising an important question. This work was primarily supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0021221 (A. L. C.). Spin-wave calculations (P. A. M.) were supported by the Russian Science Foundation via project 23-12-00159. The work by S. J. was supported by the National Science Foundation under DMR-2110041. S. J. is also supported by the Department of Energy (DOE), Office of Sciences, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract No. DEAC02-76SF00515. A. L. C. would like to thank Aspen Center for Physics and the Kavli Institute for Theoretical Physics (KITP) where different stages of this work were advanced. The Aspen Center for Physics is supported by National Science Foundation Grant No. PHY-2210452 and KITP is supported by the National Science Foundation under Grant No. NSF PHY-2309135.